OpenStudy (anonymous):

if sine(x) = 3/4, find cos(2x). --how do you solve this?

OpenStudy (anonymous):

\[\sin^{-1} \frac{ 3 }{4 } = x \] After you solve for X plug it in to cos(2x)

OpenStudy (anonymous):

Use the inverse of sin to solve for x, and then plug in (is the explanation).

OpenStudy (anonymous):

thankyou that works! :)

OpenStudy (jhannybean):

\[\large \sin(x)=\frac{3}{4} \implies \sin^{-1}(\frac{3}{4})=x\]\[\large \cos(2*(\color{red}{\sin^{-1}(\frac{3}{4})}) =?\]

OpenStudy (jhannybean):

There ya goooo. good job.